# Capsule: Efficient Player Isolation for Datacenters

### v1

@@ Abstract.
Cloud gaming is increasingly popular. A challenge for cloud provider is to keep datacenter utilization high: a non-trivial task due to application variety. These applications come in different shapes and sizes. So do cloud datacenter resources, e.g., CPUs, GPUs, NPUs.
Part of the challenge stems from game engines being predominantly designed to run only one player. One player in a lightweight game might utilize only a fraction of the cloud server GPU. The remaining GPU capacity will be left underutilized, an undesired outcome for the cloud provider. We introduce Capsule, a mechanism that allows multiple players to seamlessly share one GPU.
We implemented Capsule in O3DE, a popular open source game engine. Our evaluations show that Capsule can increase datacenter resource utilization by accommodating up to 2.25x more players, without degrading player gaming experience. Capsule is also application agnostic. We ran four applications on Capsule-based O3DE with no application changes. Our experiences show that Capsule design can be adopted by other game engines to increase datacenter utilization across cloud providers.
copyright: none
@@ 1. Introduction
However, it is challenging to achieve high datacenter utilization with gaming applications. Part of the challenge arises from games having diverse shapes and sizes. Shapes correspond to different resources games consume, such as CPU, RAM, GPU, and VRAM. Sizes correspond to their different amount, e.g., a graphics-intensive game consumes the entire GPU while a graphics-light game consumes only fraction of the GPU. Another challenge arises from datacenters having diverse hardware. A datacenter has servers with several generations of CPUs and GPUs. For example, a server with the oldest GPU could accommodate only one player while a server with the latest GPU accommodates dozens. Thus, cloud providers need a mechanism to share GPU, as well as other resources across multiple players.
Figure 2. Existing isolation mechanisms in the cloud. Lambdas are lightweight but require knowledge of application semantics. Virtual Machines are agnostic to semantics but are heavyweight. Container approach strikes the right balance for game engines.
@@
Games are not the only kind of application that require cloud resource multiplexing. Other cloud workloads also suffer this problem and resource virtualization has been the primary solution for over 20 years (Barham et al., 2003). Broadly, there are three virtualization categories: Virtual Machines (VM), containers, and lambdas, as shown in Figure 2. The container approach is well-fit for gaming applications because (1) it does not require knowing the application semantics, which makes the solution applicable to wide range of games, and (2) it does not impose high overhead, i.e., multiplexing mechanism itself consumes insignificant CPU, RAM, and GPU resources. Lambdas suffer (1) and VMs suffer (2). How does container-like multiplexing solution look like for gaming applications?
@@
We propose player-level multiplexing. A player in the graphics-heavy application will continue consuming the entire GPU. However, when a GPU has sufficient capacity to accommodate two or more players in a multiplayer game, GPU resources will be multiplexed across these players. We designed, implemented, and evaluated Capsule: an in-game-engine player isolation mechanism for multiplayer games. Capsule also allows cross-player sharing without cross-player performance inference. For example, when two players enter a room and have a shared game asset in their view, we can reuse the asset geometry across these two players, without players noticing. Sharing allows cloud providers to have a sublinear resource consumption growth for linear player increase–the phenomena we call super-linear player count, or superlinear for short.
@@
We implemented Capsule in O3DE, a popular open source game engine (O3DE, 2025b). Our evaluations show that Capsule-based O3DE can increase datacenter resource utilization by accommodating up to 2.25x more players, without degrading player gaming experience. Capsule is also application agnostic. We ran four applications on Capsule with no application changes. Our experiences show that Capsule design is generalizable and can be adopted by other engines.
@@ 2. Design and Implementation
There are four requirements for a player-isolation mechanism in cloud, in order of their priority:
  • R1: Transparent: Players should be unaware of other players sharing cloud resources. A player experience, such as, streaming quality and frames-per-second (FPS), should not degrade due to other players.
  • R2: Compatible: Do not require significant changes to run existing applications, best if no application changes. The workflow for developing a new application should also remain near identical, if not exactly identical.
  • R3: Lightweight: Isolation should not consume significant system resources, such as CPU and RAM.
  • R4: Efficient: Maximize cross-player sharing, i.e., superlinearity. For example, resource footprint (e.g., RAM) of the second player should be less than that of the first player because of sharing.
Capsule satisfies all four requirements. However, for example, the process-level-isolation would violate R1, R3, and R4. The process-level-isolation is achieved when we simply run separate game engine process in a cloud server for each player. Here, R1 is violated because a player can consume more GPU resources during graphics-heavy scene, reducing the FPS for other players. R3 is violated because spawning and managing separate OS process is not lightweight. R4 is violated because players are unable to share computation results–a by-product of process-level-isolation by design.
However, Capsule introduces player fate-sharing, which might degrade player fault-tolerance. For example, when a server GPU fails, all players running on the game engine hosted on that GPU also fail. However, fate-sharing is common and inevitable in cloud: VMs, containers, lambdas all suffer it (Figure 2), e.g., when a server fails all VMs hosted on that server fail. The disadvantages of the fate-sharing can be ameliorated by other mechanisms, such as player migration with player state replication, as it has been done for VMs over 20 years ago (Clark et al., 2005).
@@
Figure 3 shows Capsule architecture, along with other essential modules in cloud deployment. Capsule is a new module in O3DE. Capsule communicates with different system components, such as audio system, input system, rendering system (includes both audio and video rendering), and game logic (event system). Capsule leverages Entity–Component–System (ECS), an existing game engine architecture (Wikipedia, 2013). ECS makes it convenient to represent game world objects. An ECS-based game engine contains entities that have data components and systems to operate on those components. This architecture is widely adopted by modern game engines (Unity, 2025), including O3DE.
@@
Capsule distinguishes different players’ input, output, and behaviours by using Capsule Storage. Capsule Storage has two layers: entities and events. Entities are game objects, such as a car, a light, an avatar. These are the key elements that decide what the final rendered video or audio is. Events, or gameplay events, include in-game events, such as running, jumping, exploding. They are predesigned by game developers and written inside game script. These are the key elements that decide the behavior of each player and the behaviour of global environment.
@@
As shown in Figure 3, Capsule has two types of Storage: global and local. There is only one global storage in the entire engine. All players share the entities and events inside the global storage. There are one or more local Capsule Storages, one for each player. Per-player events are contained within the local storage. Capsule automatically isolates player-specific tasks at runtime by directing per-player entities and events to the respective player-owned local storage.
@@
We ported four applications to Capsule-based O3DE. Figure 1 shows three of them for brevity. Our experience with these four applications show that Capsule-based O3DE is fully compatible (R2) with non-Capsule O3DE. No changes were required on these applications to run them on Capsule-based O3DE. Moreover, the workflow for further development of these applications on Capsule-based O3DE was identical to that of non-Capsule O3DE.
@@
Capsule design principles can be easily adopted in other game engines thanks to ECS architecture. We also believe that four design requirements (R1R4) are the common goals in all cloud deployments. Thus, our design and implementation decisions are widely applicable to other game engines as well as other cloud environments.
@@ 3. Evaluation
We evaluated Capsule on a Windows workstation with AMD Ryzen 7 5800X 8-Core processor, 32 GB of RAM, and NVIDIA GeForce RTX 4090 GPU with 24 GB of VRAM. We fixed the application FPS to 30, a common minimum threshold. We compare Capsule resource footprints against the baseline that is implemented as process-level-isolation, i.e., a separate game engine process for each player. Note that this is an overly optimistic baseline for lightweightness comparison because it does not satisfy the transparency requirement (R1), as mentioned earlier. A realistic baseline that satisfies this requirement would inherently be less lightweight. Thus, the measured benefits of Capsule are conservative.
Figure 1 shows that Capsule can accommodate up to 8 players without dropping FPS below the threshold. Figure 1(a) and Figure 1(b) show game server view with single player (1p) and two players (2p), respectively. Figure 1(c) shows resource utilizations for these two environments. As shown in Figure 1(c), 2p GPU utilization (57%–67%) is up to 39% point higher than the baseline (28%–44%). This delta is smaller, but is more significant in other resources. If the delta is zero (1p and 2p lines overlap), it means 2p utilization of that resource is identical to that of 1p, which means the second player came for free (for that resource). In Figure 1(c), this is almost the case for CPU, RAM, and VRAM, i.e., the second player has a small extra cost. Note that Capsule is unable to accommodate more than two players because the FPS drops below the threshold (30) with the third players. This is not visible in Figure 1(c) because there CPU utilization is plotted for all 8 cores while there is only one main thread for all players, which runs on a single CPU core. That thread is the bottleneck. We can improve Capsule implementation to have separate main thread per player, a future work.
@@
The same conclusion holds in Figure 1(d)–(f) and Figure 1(g)–(i), for O3DE Multiplayer Sample (O3DE, 2025a) and Exhibition, respectively: the resource footprint of additional players is sublinear. For example, as shown in Figure 1(i) for Exhibition application, one player consumes 18% of GPU while 8 players consume around 99%: a sublinear per-player increase. Similarly, CPU consumption increases sublinearly: only by up to 21% point with 8 players (28%–37%) vs. one player (16%-18%). Note that O3DE Multiplayer Sample also suffers aforementioned single-thread bottleneck. Exhibition does not suffer it and therefore can achieve over 99% GPU utilization for 8 players.
@@
Figure 4. Scalability of Capsule as we increase number of players. Capsule and baseline could host up to 9 and 4 players, respectively, on the Exhibition application. Capsule accommodates up to 2.25x more players thanks to sublinear resources increase per added player.

### v3

@@ Abstract.
Cloud gaming is increasingly popular. A challenge for cloud provider is to keep datacenter utilization high: a non-trivial task due to application variety. These applications come in different shapes and sizes. So do cloud datacenter resources, e.g., CPUs, GPUs, NPUs. Part of the challenge stems from game engines being predominantly designed to run only one player. For example, one player in a lightweight game might utilize only a fraction of the cloud server GPU. The remaining GPU capacity will be left underutilized, an undesired outcome for the cloud provider.
We introduce Capsule, a mechanism to seamlessly share one GPU, and other cloud servers resources, across multiple players. Sharing makes the cost of multiple players sublinear. We implemented Capsule in O3DE, a popular open source game engine. Our evaluations show that Capsule increases datacenter resource utilization by accommodating up to 2.25x more players, without degrading player gaming experience. This is the product of Capsule using up to 1.43x less GPU, 3.11x less VRAM, 3.7x less CPU, and 3.87x less RAM compared to the baseline. Capsule is also application agnostic. We ran four applications on Capsule-based O3DE with no application changes. Our experiences with four applications, three servers with different hardware specifications, including the one with four GPUs, and multi-server cluster show that Capsule design can be adopted by other game engines to increase datacenter utilization across cloud providers.
conference: ; Under review; 2025copyright: noneccs: Computer systems organization Cloud computingccs: Computer systems organization Real-time system architectureccs: Computing methodologies Graphics systems and interfaces
@@ 1. Introduction
However, it is challenging to achieve high datacenter utilization with gaming applications. Part of the challenge arises from games having diverse shapes and sizes. Shapes correspond to diverse resources games consume, such as CUDA cores, RT cores, and Tensor cores in GPUs, in addition to the host CPU and RAM. Sizes correspond to the differing amount of these resources games consume, e.g., a graphics-intensive game consumes the entire GPU while a graphics-light game consumes only a fraction of that GPU. Another challenge arises from datacenters having diverse hardware. A datacenter has servers with several generations of CPUs and GPUs (Patel et al., 2023). For example, a server with the oldest GPU can accommodate only one player while with the latest GPU accommodates dozen players. Thus, cloud providers need a mechanism to share GPU, and other resources, across multiple players.
@@
Games are not the only kind of cloud application that require server resource multiplexing. Other cloud workloads do too and resource virtualization has been the primary solution for over 20 years (Barham et al., 2003). Broadly, there are three virtualization categories: Virtual Machines (VM) (AWS, 2025), containers (Google, 2025), and lambdas (Azure, 2025), as shown in Figure 2. The container approach is well-fit for games because (1) it does not require knowing the application semantics, which makes the solution applicable to wide range of games, and (2) it does not impose high overhead, i.e., multiplexing mechanism itself consumes insignificant resources, e.g., CPU cycles. Lambdas suffer (1) and VMs suffer (2). How does container-like multiplexing solution look like for gaming applications?
@@
We propose player-level multiplexing. A player in the graphics-heavy application will continue consuming the entire GPU. However, when a GPU has sufficient capacity to accommodate two or more players in a multiplayer game, GPU resources will be multiplexed across these players. We designed, implemented, and evaluated Capsule: an in-game-engine player isolation mechanism for multiplayer games. Capsule also allows cross-player sharing. For example, when two players enter a room and have a shared game asset in their view, we can reuse the asset geometry across these two players, without players noticing. Sharing offers cloud providers with sublinear resource usage growth for linear player increase: the phenomena we call sublinear resource footprint.
Figure 2. Existing isolation mechanisms in the cloud. Lambdas are lightweight but require knowledge of application semantics. Virtual Machines are agnostic to semantics but are heavyweight. The container approach strikes the right balance for gaming applications.
@@
We implemented Capsule in O3DE, a popular open source game engine (O3DE, 2025b). Our evaluations show that Capsule-based O3DE can increase datacenter resource utilization by accommodating up to 2.25x more players, without degrading player experience. Capsule is also application agnostic. We ran four applications on Capsule with no application changes. Our experiences show that Capsule design is generalizable and can be adopted by other engines to increase datacenter utilization across cloud providers.
2. Requirements
There are four requirements for a player-isolation mechanism in cloud, in order of their importance:
  • R1: Transparent: Players should be unaware of other players sharing cloud resources. A player experience, such as, input latency, output streaming quality, and frames-per-second (FPS), should not degrade due to other players.
  • R2: Compatible: Player isolation should not require significant changes to run existing applications, best if no application changes are required. The workflow for developing a new application should also remain near identical, if not exactly identical.
  • R3: Lightweight: Isolation mechanism itself should not consume significant system resources, e.g., CPU and RAM.
  • R4: Efficient: Maximize cross-player sharing. For example, resource (e.g., CPU) footprint of the second player should be less than that of the first player because the second player can reuse some computation results from the first player.
Capsule satisfies all four requirements. However, for example, the process-level-isolation would satisfy R1 and R2, but violate R3 and R4. The process-level-isolation is achieved when we simply run separate game engine process in a cloud server for each player. R3 is violated because spawning and managing separate OS process is not as lightweight as handling all computation within the same process. R4 is violated because players are unable to share computation results, which is the by-product of the process-level-isolation by design because the OS processes operate on a separate memory address (unless another mechanism is used for cross-process-memory-sharing). See Section 6 for further discussion of functional and performance transparency (R1).
@@ 3. Design and Implementation
We designed Capsule to satisfy all four aforementioned requirements. Figure 3 shows Capsule architecture, along with other essential modules in cloud deployment. Capsule is a new module in O3DE. Capsule communicates with different system components, such as audio system, input system, rendering system (includes both audio and video rendering), and game logic (event system). Capsule leverages Entity–Component–System (ECS), an existing game engine architecture (Wikipedia, 2013). ECS makes it convenient to represent game world objects. An ECS-based game engine contains entities that have data components and systems to operate on those components. This architecture is widely adopted by modern game engines (Unity, 2025), including O3DE.
As shown in Figure 3, players connect to the cloud over the wide area network, e.g., Internet. Players’ entry point is the Streaming module, which creates a separate game session for each player, isolating their inputs, such as keyboard and mouse, and outputs, such as video streams. The Streaming module passes the player-specific input to a separate game session, which includes player-specific UI, Rendering, Input, Audio, and Encoding states. These states are reflected in the game, but are managed inside the engine. In other words, application will perceive all player-specific states to be isolated from each other at the engine level, i.e., as if each player (or in fact, each game instance) has an engine of its own. In Capsule-based O3DE, all of these players share one engine.
@@
Capsule distinguishes different players’ input, output, and behaviours by using Capsule Storage. Capsule Storage manages two constructs: entities and events. Entities are game objects, such as a car, a light, an avatar. Events, or gameplay events, include in-game events, such as running, jumping, exploding. They are predesigned by the game developer and are written in the game script. Entities and events determine the behavior of each player and of the global environment. They also determine the final rendered frame.
@@
As shown in Figure 3, Capsule has two types of Storage: global and local. There is only one global storage in the entire engine. All players share the entities and events inside the global storage. There are one or more local Capsule Storages, one for each player. Entities inside the local storage are visible only to storage-owner-player except for the player controlled network entities, such as the player avatar whose state is shared with other players. For example, when a player jumps, that avatar jump should be visible to everyone. Per-player events are contained within the local storage. Capsule isolates player-specific tasks at runtime by directing player-specific entities and events to the respective player-owned local storage.
Figure 4 shows how Capsule handles per-player entity tracking using global and local storage. There are seven kinds of constructs, starting with Predefined Global Entity and ending with Player Controlled Network Entity. Not all of these constructs are exposed to the game developer. In fact, the game developers design games as if they did for vanilla (non-Capsule) O3DE, e.g., declare an entity as Network Entity (such as a museum piece in Exhibition) or Player Controlled Network Entity (such as a player avatar). Capsule manages entity division into local and global, as well as reference tracking. For example, when there are no players, there are no constructs. Optionally, Capsule could automatically create and maintain a game level cache to accelerate game loading when players join. This cache is invisible to the game.
As Figure 4 shows, players join through API calls from the Streaming module (see Figure 3). When the API call is made for the first player, predefined global entities are initialized, unless they have already been initialized via cache. During player join, all entities that the game developer specified in the initial level are spawned. For subsequent players, predefined global entities are not re-instantiated, instead, references to the original entities are created within each subsequent player’s local storage. These references ensure that the player local gameplay logic remains consistent with the vanilla game-level design (for non-Capsule engine).
When the second player joins, local entities are duplicated for each player. A corresponding reference for each duplicate is maintained in the global storage. This reference mechanism ensures that player-specific gameplay logic is consistent with the vanilla design. Network entities, i.e., entities whose state and logic must be synchronized through the game server (Lu et al., 2006), are typically treated as global entities since network entities are inherently shared among players. Such treatment is consistent with legacy multiplayer architecture.
Event tracking follows analogous flow as the entity tracking. There are global events, such as time-of-the-day change in Paris Opera House, and local events, such as bullet inventory decrease after player shots in O3DE Multiplayer Sample. These distinctions also apply to the event storage, which also separates global and local scopes. Global events may trigger both global and local events, whereas local events are restricted to interactions within the same local event storage. For example, time-of-the-day change event affects all players, whereas a local inventory-update event should remain confined to that player’s local storage.
A notable exception arises with player-controlled network entities, e.g., the entity that is labelled with “Dynamically Added from Game Server” in Figure 4. This entity is globally visible across all players, thereby functioning as a global entity, but its event propagation is restricted to local event storage. For instance, a player pressing the “attack” command generates its own animations and state changes without broadcasting to other players. However, when this action results in interactions with other players (e.g., a collision), the corresponding events are escalated to the game server and synchronized globally, making these events visible to other players. This local-propagation-until-interaction approach ensures that player inputs remain isolated, while cross-player interactions get broadcasted across the entire shared game world.
Figure 4. Player entity tracking using global and local Capsule storage. Global storage is instantiated only once. Local storage is created as players join and is destroyed when they leave the game.
We designed Capsule Storage (Figure 4) and Capsule architecture (Figure 3) with generality in mind. Most of the Capsule changes lies on either the ECS layer or the input and output systems of the game engine. This design maintains compatibility with other sub-systems, such as the rendering and physics systems. Therefore, it is possible, and in fact, it is encouraged, to replace these systems with multi-player-aware alternatives, without modifying Capsule. For example, a related work, On Surface Caches (OSC) (Weinrauch et al., 2023), caches and reuses the computed color values of the same world position across multiple players. The default O3DE rendering sub-system can be replaced with OSC-like alternative to further improve cross-player-sharing, without changing Capsule.
@@
We ported four applications to Capsule-based O3DE to validate our design. All four applications nicely fit within our entity and event tracking systems. Figure 1 shows three of them for brevity. Our experience with these four applications show that Capsule-based O3DE is fully compatible (R2) with non-Capsule O3DE. No changes were required to these applications. Moreover, the workflow for further development of these applications on Capsule-based O3DE was identical to that of non-Capsule O3DE.
@@
Capsule design principles can be easily adopted in other game engines thanks to ECS architecture. We also believe that four design requirements (R1R4) are the common goals in all cloud deployments. Thus, our design and implementation decisions are applicable to other game engines and other cloud environments.
4. Methodology
Table 1. Workstations used for Capsule evaluation. All workstations use NVIDIA GPUs. RAM and VRAM values are in GB.
Workstation CPU CPU Cores RAM GPU VRAM GPU Count
SingleGPU AMD Ryzen 7 5800X 8 32 GeForce RTX 4090 24 1
DualGPU 13th Gen Intel Core i7-13700K 16 64 GeForce RTX 3090 24 2
QuadGPU AMD Ryzen Threadripper PRO 5975WX 32 256 Fortis  [^1] Exact GPU model is not disclosed. VRAM “24+” means it has more VRAM than the above two. Fortis is our custom label, means strong in Latin. 24+ 4
We evaluated Capsule on diverse datacenter hardware. We used three different workstations: with one GPU, with two GPUs, and with four GPUs, as described in Table 1. All workstations run Windows OS to faithfully reproduce our production environment. Capsule implementation exists for Linux OS but is not as thoroughly performance evaluated as Windows. We believe results we report here are applicable to Linux environment as well. We fixed the application FPS to 30, a common minimum threshold. We read the system-wide utilization levels of GPU, VRAM, CPU, and RAM resources every second. The value on each second is the average utilization during that one-second interval; consistent with Windows performance counters (Microsoft, 2025).
We compare the cloud server resource consumption of Capsule against the Baseline. For Baseline, we implemented process-level-isolation, i.e., a separate game engine process for each player. (We were unable to use production-level alternative virtualization techniques, such as NVIDIA RTX Virtual Workstation (vWS) (NVIDIA, 2025b), for the baseline due to NVIDIA licensing restrictions (NVIDIA, 2025c).) In Baseline, we launched the Exhibition game server (Lu et al., 2006) and then launched game clients one by one, measuring the server utilizations as players get added. The client-side evaluation is identical between Capsule and Baseline, but on the server side, after the first player, we keep adding players to the same client process (running in the cloud server), rather than creating a separate process per player (in the cloud server). This is consistent with the Capsule design in Figure 3.
We evaluated Capsule’s applicability to diverse datacenter workloads by running three different applications–Paris Opera House, O3DE Multiplayer Sample (O3DE, 2025a), Exhibition–on SingleGPU workstation (see Table 1). We monitored resource utilization for 40 seconds of gameplay time.
In hardware diversity experiments, we ran Exhibition on three workstations from Table 1. In these experiments, we compare resource utilizations of Capsule and Baseline with two or more players. In each experiment, the gameplay of each player is unique, and is deterministic across different hardware. It is unique because the player walking trajectories differ within the same experiment. We record a stochastic trajectory for each player in a separate experiment and replay that trajectory in Capsule vs. Baseline evaluations. For example, if there are 4 players in the experiment, each player trajectory is stochastic. Stochasity makes our evaluations unbiased, i.e., free from the selective benchmarking crime (Heiser, 2025). The view the player gets in a frame influences the rendering load for that frame, which in turn influences how much computation can be shared across multiple players in that frame. In the biased case, we would have all players have the same view, or largely overlapping view, which would unfairly make Capsule outshine the Baseline because the amount of cross-player sharing is maximized. By adopting stochasity, our evaluations are free from such bias. In fact, stochastic trajectories might undersell the Capsule benefits because players might inadvertently have less view overlap, hence less sharing, than they would otherwise have in the realistic, production deployment. We would rather undersell than be biased. Thus, the Capsule benefits we report in our experiments are conservative.
The determinism across different hardware allows our experiments stay true to our claim: hardware is different, everything else is the same. When we evaluate different hardware, we replay the prerecorded trajectories for the first player, the second player, and so on. Thus, the player generated rendering load is identical across all hardware configurations while the number of players accommodated might increase (or decrease) depending on the compute capacity of the hardware under evaluation. This exactly is the purpose of the hardware diversity experiments: evaluate if the Capsule benefits are consistent across different datacenter hardware.
In general, we chose the strongest viable baseline and compared it to Capsule. We evaluated on wide range of applications and diverse datacenter hardware while faithfully replicating our production environment and remaining bias free.
@@ 5. Evaluation
We evaluate how efficiently Capsule accommodates multiple players. We run three applications on three datacenter servers. We first evaluate all three applications on SingleGPU workstation, with the strongest GPU. We then evaluate Exhibition application on three workstations, and on a multi-server cluster.
Figure 1 shows three applications on SingleGPU workstation. Figure 1(a) and Figure 1(b) show game server view of Paris Opera House with single player (1p) and two players (2p), respectively. Figure 1(c) shows resource utilizations for these two environments. Unlike other applications used in our experiments, GPU usage in Paris Opera House has high variance. For example, in some small intervals, 2p utilization line is below 1p utilization. This is partially due to the utilization capturing noise, but is mostly due to view angle of the players. For example, between 25th-28th seconds, if two players in 2p environment happen to view the corner of the wall, while the single player in 1p environment views the distant place with many polygons, 1p GPU utilization will be higher. However, if we summarize the GPU utilizations across the entire 40-second gameplay, without focusing on a small time intervals, the average 2p GPU utilization ($\approx$50%) is $\approx$10% higher than 1p ($\approx$40%). This is expected: with Capsule, the second player imposes a sublinear $\approx$10% GPU cost.
This delta is smaller, but is more significant in other resources (VRAM, CPU, RAM). If the delta is the smallest, i.e., zero, 1p and 2p lines overlap. This means that 2p utilization of that resource is identical to that of 1p, which means the second player came for free (for that resource). In Figure 1(c), this is almost the case for VRAM, CPU, and RAM: the second player has a small extra cost. Thus, the cost of additional players is sublinear for these resources.
Capsule is unable to accommodate more than two players in Paris Opera House application. The FPS drops below the threshold (30) with the third player due to CPU bottleneck. This is not visible in Figure 1(c), i.e., CPU utilization is only $\approx$30% for 1p as well as 2p. This is misleading because CPU utilizations are reported and are plotted for all 8 cores (see SingleGPU workstation in Table 1) while there is only one main thread for all players, which runs on a single CPU core. That thread is the bottleneck. We can improve Capsule performance by spreading players across CPU cores. We have not done so, yet, because in most games, GPU is the bottleneck (as evident in our third application, Exhibition). However, extending Capsule to use different CPU cores for different players is the right future work to make Capsule applicable to diverse applications.
@@
The sublinear cost conclusion holds in O3DE Multiplayer Sample (O3DE, 2025a) and Exhibition applications, in Figure 1(d)–(f) and Figure 1(g)–(i), respectively. For example, as shown in Figure 1(i) for Exhibition application, one player consumes $\approx$18% of GPU while 9 players consume around $\approx$99%: a sublinear per-player increase. Similarly, CPU consumption increases sublinearly: only by $\approx$10% point with 9 players ($\approx$29%) vs. one player ($\approx$19%). Note that O3DE Multiplayer Sample also suffers aforementioned single-thread bottleneck. Exhibition does not suffer it and therefore can achieve over 99% GPU utilization for 9 players. These results demonstrate that Capsule, as it is implemented now, brings greater benefit to graphics-heavy application, i.e., when GPU is bottleneck. Our future work will alleviate CPU bottleneck.
@@
Figure 5. Scalability of Capsule as we increase number of players. Capsule and Baseline host up to 9 and 4 players, respectively, on the Exhibition application. Capsule accommodates up to 2.25x more players thanks to sublinear resources increase per added player.
Figure 6. System-wide utilizations with Capsule and Baseline on SingleGPU workstation.